# Compact: Common-token Optimized Model Pruning Across Channels and Tokens

### v1

@@ 4.3 Efficiency in Pruning Time, Inference Latency, and Memory
Table 4: Pruning time ($m$:$s$) comparison at a 35% pruning ratio of LLaMA 3.1.
@@
Pruning time.
The main strength of training-free pruning methods is that they are able to prune large models in a short amount of time. To measure this, we report how long our method takes to prune models at a 35% pruning ratio compared to our baselines. We report our results on LLaMA 3.1-8B and 70B, as larger models enable us to better discriminate between different methods. For fairness, we exclude the time taken to load and save the models and data. All reported results are an average of 3 different pruning runs.
Our results are in Table 4. With LLaMA 3.1-8B, Compact is 3 times faster than 2SSP, our strongest baseline, and within an order of magnitude of our depth pruning methods, pruning the model in less than a minute. At the 70B size, Compact now becomes 6 times faster than 2SSP, and the gap between ShortGPT and Compact shrinks to under 10 seconds. The low pruning times show that Compact has comparable efficiency to our baselines.
@@
Settings: i) Text classification: When running our pruned models on our downstream performance benchmarks, we record the maximum memory usage as well as the throughput, measured in number of questions per second. For this test, we use the HellaSwag benchmark, as it is the longest test in our benchmark suite, which allows us to better discriminate between the methods. For the same reason, we test on LLaMA 3.1-8B using a 3-run average. All models are loaded in 16-bit precision on a single A100-80GB GPU, with a batch size of 256. ii) Text generation: We use the vLLM library (Kwon et al., 2023) to test the memory reduction and inference speedup of our method. As mentioned before, because SliceGPT and 2SSP are incompatible with vLLM, we omit it from our tests. Similarly to the text classification test, we use a 3-run average of LLaMA 3.1-8B in 16-bit precision on a single A100-80GB GPU, but with a batch size of 1 instead, as this is a more realistic workload for text generation. Tests are conducted with 128 input tokens and 128 output tokens.
Our results are in Table 5. Note that ShortGPT and LaCo prune the same number of layers, so the resulting pruned model has an identical architecture. As a result, they have the same inference performance. In the text classification task, our method achieves the highest memory reduction, and the second-highest throughput increase. We attribute the low memory usage to the reduced vocabulary size. During the forward pass, the model logits are stored in GPU memory, which has size $\text{Batch Size}\times\text{Seq Len}\times\text{Vocab Size}$. This matrix can become very large in scenarios with high batch sizes, reaching the tens of gigabytes in this test. By pruning the vocabulary size, Compact significantly reduces the memory usage over other pruning methods. Compact has higher throughput than our width pruning baselines SliceGPT and 2SSP, although it falls behind our depth pruning methods. This mirrors findings in Bian et al. (2025), which showed that scaling down the number of layers results in a proportional increase in throughput, but scaling down the size of each layer does not. This is because matrix multiplications in each layer are performed in parallel, and reducing the matrix size only results in lower GPU utilization with a minimal throughput increase. These results suggest that while Compact achieves faster inference compared to other width pruning methods, further work is needed to match the inference speed of depth pruning methods. In the generation task, the trends in memory usage and throughput are similar to that of the classification task.
@@ 4.4 Recovery Fine-tuning
Although Compact is a training-free method, pruned models can be fine-tuned afterwards to recover performance. Due to the page limit, our recovery-finetuning experiments can be found at A.4.
@@ 5 Ablation Study
In ablation study, we try to further three questions: i) $\mathcal{Q}_{1}$: how effective our intermediate pruning method Common-act2? ii) $\mathcal{Q}_{2}$: How much the various configuration of Vocab-pruning and FFN-pruning lead to? iii) $\mathcal{Q}_{3}$: how much calibration data is needed?
@@
Figure 2: Comparison of different intermediate pruning methods on Qwen 2.5-0.5B at a 35% pruning ratio when combined with rare vocabulary removal. We see that common act2 outperforms both act2 and ||act||.
@@
Compact has two hyperparameters: the new vocabulary size $V^{\prime}$ and the new intermediate size $I^{\prime}$. These hyperparameters can be adjusted to produce many different model configurations at a given pruning ratio. We test different configurations of Qwen 2.5-0.5B at a 35% pruning ratio by starting with $V^{\prime}=151936,I^{\prime}=2048$. This configuration leaves $V^{\prime}$ unchanged, which is identical to act2. Then, we progressively increase $I^{\prime}$ and decrease $V^{\prime}$ until $V^{\prime}$ reaches 0. This results in 8 different model configurations to test.
The results can be seen in Table 6. Despite act2’s simplicity, it already achieves better performance than our baselines on its own, without any rare token removal. However, we find that the best result is achieved with a combination of Vocab-pruning and FFN-pruning, validating the effectiveness of our proposed method.
@@
V’ I’ Para (M) Ratio Avg Avg% Dense 151936 4864 494 0.00% 50.2 100.0 act2 151936 2048 312 36.84% 31.6 63.0 Compact 131584 2304 311 37.04% 31.8 63.3 111104 2560 309 37.45% 33.1 66.1 90752 2944 315 36.23% 34.2 68.2 70400 3200 314 36.44% 34.5 68.9 49536 3456 311 37.04% 35.3 70.4 29568 3712 310 37.25% 34.7 69.2 9088 3968 308 37.65% 32.4 64.7
@@
Figure 3: Downstream performance across different calibration dataset sizes. Although all benchmarks used 256 calibration samples, 16 samples gives similar performance.
816326412825651210246666686870707272# SamplesAvg%
All of our experiments were conducted using 256 calibration samples from the C4 dataset. We perform ablations over the number of calibration samples used, with our results in Figure 3. We find that Compact is highly robust to the number of calibration samples, and similar performance can be achieved with just 16 samples, implying that the pruning time can be reduced even further.
@@ 6 Conclusion
We proposed Compact, a training-free pruning method combining vocabulary and FFN pruning under the common-token distribution. Experiments show that it achieves robust performance across model scales, offering smooth degradation, strong efficiency, and broad deployment compatibility.
@@ A.1 Parameter Distribution Across Other Model Families
Figures 5 and 5 provide parameter distributions for the LLaMA 3 and Gemma 3 model families, respectively. We see that these models follow the same trend as Qwen 2.5 where smaller LLMs have a higher proportion of vocabulary parameters, corroborating our theoretical analysis.
@@ A.2 Evaluation Methodology
Downstream evaluations were conducted using the LM-Evaluation-Harness (Gao et al., 2024a). The evaluation details for each benchmark are in Table 7.
@@ A.3 Pruning Hyperparameters
A.4 Recovery Fine-tuning
Although Compact is a training-free method, we can optionally perform recovery fine-tuning after pruning to regain additional performance. To test this, we perform full finetuning on Qwen 2.5-0.5B after pruning at a 35% pruning ratio. We conduct two experiments with this setup:
  1. 1.
    Continued Pretraining (CPT): We use the first 900 million tokens of the FineWeb-Edu dataset (Penedo et al., 2024), a high quality pretraining dataset that achieves SOTA performance compared to other datasets using an equal number of tokens.
  2. 2.
    Self-Data Distillation (SDD): Modern model families like Qwen use very high-quality, private pretraining datasets that have a different distribution than open-source versions. As a result, performing continued pretraining could result in worse performance. We use a modified version of self-data distillation (Thangarasa et al., 2024), which creates a dataset with the same distribution as the training dataset.
    More specifically, we use the unpruned model to generate 900 million tokens of text completions. These completions require at least one input token to begin generation, so we use the first token of examples in FineWeb-Edu. We generate these completions with temperature=1.0 and no other sampling parameters to ensure that the dataset reflects the unmodified probability distribution of the unpruned model. After this, we fine-tune our pruned model on this synthetic dataset.
Our results are shown in Table 9. By manually examining samples from the synthetic dataset, we find that many of the completions are incoherent, often degenerating into repetitive text and hallucinations. This is due to using a tiny teacher model in conjunction with our lack of sampling parameters. Despite this, we find that self-data-distillation significantly outperforms CPT, achieving 5% higher performance recovery and 13% total performance recovery compared to the training-free method. Although CPT achieves higher scores in a majority of the benchmarks, it does not achieve any improvements in MMLU or GSM8K, our two hardest benchmarks. In contrast, SDD is able to consistently improve on all benchmarks, enabling a higher average score. These results highlight that SDD outperforms regular fine-tuning even at the sub-1B parameter range, extending the results in Thangarasa et al. (2024). Notably, our pruned model fine-tuned on 900 million tokens outperforms the recently-released Gemma 3-270M, despite the latter being pretrained on 6 trillion tokens. These results highlight the practicality of pruning techniques compared to pretraining.
Table 9: Comparison of post-pruning recovery fine-tuning methods.
Method Params (M) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% Dense 494 47.3 52.2 56.4 32.3 58.2 69.9 34.9 50.2 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 57.0 Ours 311 25.5 40.0 53.8 25.2 40.0 62.2 0.5 35.3 70.4 + SDD 311 39.7 43.0 54.9 28.6 50.6 65.6 11.3 41.9 83.6 + CPT 311 25.8 45.1 55.8 27.9 52.7 67.7 0.0 39.3 78.3 Gemma 3-270M 268 26.2 41.5 53.8 28.4 57.2 68.4 1.2 39.5 -

### v2

@@ 4.3 Efficiency in Pruning Time, Inference Latency, and Memory
Pruning time.
The main strength of training-free methods is that they can prune large models efficiently. To test this, we report pruning times at 35% pruning on LLaMA 3.1-8B and 70B to best discriminate between methods.
Table 5: 3-run average pruning time ($m$:$s$) comparison at a 35% pruning ratio for LLaMA 3.1. We exclude I/O time for fairness.
@@
Our results are in Table 5. With LLaMA 3.1-8B, Compact is 3 times faster than 2SSP, our strongest baseline, and comparable our depth pruning methods, with pruning times under a minute. At the 70B size, Compact now becomes 6 times faster than 2SSP. The low pruning times show that Compact has competitive efficiency to our baselines.
@@
We evaluate two inference paradigms: Text classification and Text generation. The inference setup is in Appendix A.4.
Our results are in Table 6. Note that ShortGPT and LaCo prune the same number of layers, so they have the same inference performance. In the text classification task, our method achieves the highest memory reduction, and the second-highest throughput increase. The low memory usage is from the reduced vocabulary size. During the forward pass, logits are stored in GPU memory, which becomes very large with high batch sizes. By pruning the vocabulary size, Compact shrinks logit size, causing large memory reductions. This is especially important in edge computing applications, where memory is very limited and can spell the difference in whether a model can be used or not. Compact has higher throughput than our width pruning baselines SliceGPT/2SSP, although it falls behind depth pruning methods. This aligns with Bian et al. (2025), which showed that scaling down layer count proportionally increases throughput, but scaling layer size does not. While Compact achieves faster inference than other width pruning methods, more work is needed to match the throughput of depth pruning methods. In the generation task, the trends in memory usage and throughput are similar to that of the classification task.
@@ 4.4 Recovery Fine-tuning
Although Compact is training-free, we optionally apply recovery fine-tuning. We fine-tune Qwen 2.5–0.5B pruned at 35% using two approaches: (i) Continued Pretraining (CPT): train on 900M tokens from FineWeb-Edu (Penedo et al., 2024); (ii) Self-Data Distillation (SDD): generate 900M tokens from the unpruned model (temperature = 1.0) to match its training distribution, then fine-tune on this synthetic data (Thangarasa et al., 2024).
Despite some incoherence in the SDD synthetic dataset, our results (Table 7) show that SDD consistently improves all benchmarks and yields a higher average score than CPT, which fails to boost MMLU or GSM8K. This confirms SDD’s effectiveness even below 1B parameters (Thangarasa et al., 2024). With fine-tuning, Compact surpasses training-based LLM-Streamline (Chen et al., 2025) and even outperforms Gemma 3-270M, pretrained on 6T tokens—highlighting pruning’s efficiency relative to pretraining.
Table 7: Comparison of post-pruning recovery fine-tuning methods.
Method Params (M) MMLU HeSw WiGr ARC-C ARC-E PIQA GSM8K Avg Avg% Dense 494 47.3 52.2 56.4 32.3 58.2 69.9 34.9 50.2 100.0 Random - 25.0 25.0 50.0 25.0 25.0 50.0 0.0 28.6 57.0 LLM-Streamline 315 23.0 31.8 53.4 23.5 37.9 59.7 0.2 32.8 65.4 Gemma 3-270M 268 26.2 41.5 53.8 28.4 57.2 68.4 1.2 39.5 - Our Compact 311 25.5 40.0 53.8 25.2 40.0 62.2 0.5 35.3 70.4 + CPT 311 25.8 45.1 55.8 27.9 52.7 67.7 0.0 39.3 78.3 + SDD 311 39.7 43.0 54.9 28.6 50.6 65.6 11.3 41.9 83.6
Figure 2: Common-act2 outperforms both act2 and ||act|| (Qwen 2.5-0.5B at a 35% pruning ratio).
@@ 5 Ablation Study
In our ablation study, we try to further four questions: i) $\mathcal{Q}_{1}$: how effective is Common-act2? ii) $\mathcal{Q}_{2}$: how to trade off Vocab-pruning and FFN-pruning? iii) $\mathcal{Q}_{3}$: how much calibration data?
@@
816326412825651210246666686870707272# SamplesAvg%
Figure 3: Downstream performance by calibration dataset size. Although all benchmarks used 256 calibration samples, 16 samples is sufficient.
@@
Compact has two hyperparameters: the new vocabulary size $V^{\prime}$ and the new intermediate size $I^{\prime}$. These can be adjusted to produce many configurations at a given pruning ratio. We test different configurations of Qwen 2.5-0.5B at 35% pruning. Our results are in Table 8. Note that $V^{\prime}=151936,I^{\prime}=2048$ is identical to act2. Despite act2’s simplicity, it achieves better performance than our baselines even without Vocab-pruning. However, the best result is achieved with a combination of both, validating Compact’s methodology.
@@
V’ I’ Ratio Avg Avg% Dense 151936 4864 0.00% 50.2 100.0 act2 151936 2048 36.84% 31.6 63.0 Compact 131584 2304 37.04% 31.8 63.3 111104 2560 37.45% 33.1 66.1 90752 2944 36.23% 34.2 68.2 70400 3200 36.44% 34.5 68.9 49536 3456 37.04% 35.3 70.4 29568 3712 37.25% 34.7 69.2 9088 3968 37.65% 32.4 64.7
@@
We perform ablations over the number of calibration samples, with our results in Figure 3. Compact is highly robust to sample count, and similar performance can be achieved with just 16 samples, implying that the pruning time can be reduced further.
@@ 6 Conclusion
We propose Compact, a training-free pruning method combining vocabulary and FFN pruning under the common-token distribution. Experiments show that it achieves robust performance across model scales, offering smooth degradation, strong efficiency, and broad deployment compatibility. In future works, we plan to address the discrepancy in throughput between our method and ShortGPT/LaCo, closing the gap between width and depth pruning.
@@ A.1 Parameter Distribution Across Other Model Families
Figures 5 and 5 provide parameter distributions for the LLaMA 3 and Gemma 3 model families, respectively. We see that these models follow the same trend as Qwen 2.5 where SLMs have a higher proportion of vocabulary parameters, corroborating our theoretical analysis.
@@ A.2 Evaluation Methodology
Downstream evaluations were conducted using the LM-Evaluation-Harness (Gao et al., 2024a), specifically lm-eval 0.4.8. The evaluation details for each benchmark are in Table 9.
@@ A.3 Pruning Hyperparameters
A.4 Inference Settings
Settings: i) Text classification: When running our pruned models on our downstream performance benchmarks, we record the maximum memory usage as well as the throughput, measured in number of questions per second. For this test, we use the HellaSwag benchmark, as it is the longest test in our benchmark suite, which allows us to better discriminate between the methods. We test on a larger model (LLaMA 3.1-8B) using a 3-run average, again to better discriminate between methods. All models are loaded in 16-bit precision on a single A100-80GB GPU, with a batch size of 256. ii) Text generation: We use the vLLM library (Kwon et al., 2023) to test the memory reduction and inference speedup of our method. As mentioned before, because SliceGPT and 2SSP are incompatible with vLLM, we omit it from our tests. Similarly to the text classification test, we use a 3-run average of LLaMA 3.1-8B in 16-bit precision on a single A100-80GB GPU, but with a batch size of 1 instead, as this is a more realistic workload for on-device text generation. Tests are conducted with 128 input tokens and 128 output tokens.